# GraphRCG: Self-conditioned Graph Generation via Bootstrapped Representations

### v1

@@ 3.2. Self-conditioned Guidance (Training)
(7) $(\widetilde{\mathbf{p}}^{\mathbf{X}},\widetilde{\mathbf{p}}^{\mathbf{E}})=g_{\theta}(\mathbf{X}_{t},\mathbf{E}_{t},\widetilde{\mathbf{h}}_{t},t),$
where $\widetilde{\mathbf{h}}_{t}=f_{\gamma}(\mathbf{h}_{t},t)$ is the denoised representation, generated from the representation generator at timestep $t$. Moreover, $\widetilde{\mathbf{p}}^{\mathbf{E}}\in\mathbb{R}^{b}$ are the predicted distributions for the node types and the edge types, respectively. The graph generator is based on the message-passing transformer architecture (Shi et al., 2020), as it effectively extracts the complex correlations between nodes and edges, while also being suitable for incorporating representations for conditioning (Vignac et al., 2022). Specifically, the layers incorporate the graph attention mechanism (Veličković et al., 2018) into a Transformer framework (Vaswani et al., 2017), achieved by including normalization and feedforward layers. For each given (noisy) graph, the graph generator projects nodes and edges separately into low-dimensional representations and processes them through totally $L$ transformer layers, denoted as $(\mathbf{x}^{(l+1)}_{t},\mathbf{e}^{(l+1)}_{t})=M_{(l)}(\mathbf{x}^{(l+1)}_{t},\mathbf{e}^{(l+1)}_{t})$. Here $\mathbf{x}_{t}^{(l)}$ (or $\mathbf{e}_{t}^{(l)}$) is the representation for the nodes (or edges) in the $l$-th layer of the transformer, denoted as $M_{(l)}$. Based on the transformer architecture, we hereby explore several possible methods to incorporate the representation into the transformer architecture.
Figure 3. The sampling process of our framework with self-conditioned guidance. Our step-wise incorporation strategy employs the denoised presentation at each timestep to guide the denoising of the noisy graph at the same time step, thereby progressively guiding each step closer to the learned distributions.
Cross-Attention. To effectively utilize the representations for guidance, we map them to the intermediate layers of the graph transformer via the cross-attention mechanism (Vaswani et al., 2017), achieved by $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}(QK^{T}/\sqrt{d})\cdot V.$ Specifically,
(8) $Q=W_{Q}^{(l)}\cdot\mathbf{x}^{(l)},K=W_{K}^{(l)}\cdot\widetilde{\mathbf{h}}_{t},V=W_{V}^{(l+1)}\cdot\widetilde{\mathbf{h}}_{t},$
where $W_{Q}^{(l)}\in\mathbb{R}^{d\times d_{x}}$,$W_{K}^{(l)}\in\mathbb{R}^{d\times d_{h}}$, $W_{K}^{(l)}\in\mathbb{R}^{d\times d_{h}}$ are learneable projection matrices. $d_{x}$ and $d_{h}$ are the dimensions of $\mathbf{x}$ and $\mathbf{h}$, respectively. Notably, the above process is performed for edge representations $\mathbf{e}^{(l)}$ in the same way with different parameters.
@@
(9) $\mathcal{L}_{GG}=\mathbb{E}_{G,t}\left[\sum\limits_{i=1}^{n}\text{CE}(\mathbf{X}_{i},\widetilde{\mathbf{p}}^{\mathbf{X}}_{i})+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\text{CE}(\mathbf{E}_{i,j},\widetilde{\mathbf{p}}^{\mathbf{E}}_{i,j})\right],$
@@
(10) $\mathcal{L}_{AG}=\mathbb{E}_{G,t}\left[\left\|h_{\eta}(p^{\mathbf{X}},p^{\mathbf{E}})-f_{\gamma}\left(\mathbf{h}_{t},t\right)\right\|_{2}^{2}\right].$
@@
Note that $\mathcal{L}_{AG}$, together with $\mathcal{L}_{GG}$, will be only used for optimizing the graph generator, not involving the representation generator. In this way, we can ensure that the representation generator focuses on capturing the graph distribution. The overall process of self-conditioned guidance is presented in Algorithm 1.
@@
3.3. Self-conditioned Guidance (Sampling)
(11) $G_{t-1}\sim p_{\theta}(G_{t-1}|G_{t},\mathbf{h}_{t-1}),\ \ \text{where}\ \ \mathbf{h}_{t-1}=f_{\gamma}(\mathbf{h}_{t},t).$
In concrete, we keep track of the representation sampling process in the representation generator and utilize the representation in each step to guide the graph sampling in the same timestep $t$.
Table 1. Comparison of generation results on the SBM, Planar, and Ego datasets. The best results are shown in bold.
SBM Planar Ego
Training Set 0.0008 0.0332 0.0255 0.0063 0.0002 0.0310 0.0005 0.0052 0.0002 0.0100 0.0120 0.0014
SPECTRE 0.0015 0.0521 0.0412 0.0056 0.0005 0.0785 0.0012 0.0112 - - - -
HiGen 0.0019 0.0498 0.0352 0.0046 - - - - 0.0472 0.0031 0.0387 0.0062
4. Experiments
In our experiments, we evaluate our framework GraphRCG across graph datasets covering realistic molecular and synthetic non-molecular datasets. In particular, we compare our framework with other baselines, including auto-regressive models: GRAN (Liao et al., 2019) and GraphRNN (You et al., 2018b), a GAN-based model: SPECTRE (Martinkus et al., 2022), and diffusion models: EDP-GNN (Niu et al., 2020), DiGress (Vignac et al., 2022), GDSS (Jo et al., 2022), GraphARM (Kong et al., 2023), HiGen (Karami, 2023), and SparseDiff (Qin et al., 2023). We provide implementation details and hyperparameter settings in Appendix C.1.
4.1. Evaluation Metrics
For generic graph generation, we evaluate the quality of the generated graphs with structure-based evaluation metrics. We follow previous work (You et al., 2018b) and calculate the MMD (Maximum Mean Discrepancy) between the graphs in the test set and the generated graphs, regarding (1) degree distributions, (2) clustering coefficients distributions, (3) the number of orbits with four nodes, and (4) the spectra of the graphs obtained from the eigenvalues of the normalized graph Laplacian (You et al., 2018b; Chen et al., 2023). For molecular graph generation, following previous works (Jo et al., 2022; Kong et al., 2023), we evaluate the generated molecular graphs with several key metrics: (1) Frechet ChemNet Distance (FCD) (Preuer et al., 2018), which quantifies the discrepancy between the distributions of training and generated graphs, based on the activation values obtained from the penultimate layer in ChemNet. (2) Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa and De Grave, 2010), which measures the Maximum Mean Discrepancy (MMD) between the generated and the test molecules, considering both node and edge features. (3) Validity, which refers to the proportion of generated molecules that are structurally valid without requiring valency adjustments. (4) Uniqueness, which measures the proportion of unique molecules that are distinct from each other.
4.2. Generic Graph Generation
In this subsection, we further evaluate our framework on generic graph datasets with relatively larger sizes than molecular graphs. In particular, we consider two synthetic datasets: SBM, drawn from stochastic block models (Martinkus et al., 2022), with a maximum size of 200 nodes, and Planar, containing planar graphs with a fixed size of 64 (Vignac et al., 2022). In addition, we consider a realistic citation dataset Ego (Sen et al., 2008), originated from Citeseer (Giles et al., 1998). Further details of these datasets are provided in Appendix C. From the results presented in Table 1, we could obtain the following observations: (1) Our framework GraphRCG outperforms other baselines on all three datasets across various metrics for graph generation, demonstrating the effectiveness of our framework in precisely capturing graph distributions and utilizing them for generation guidance. (2) The performance improvement over other methods is more substantial on the Planar dataset. As illustrated in the t-SNE plot in Fig. 5 (b), the distribution of the Planar dataset is more scattered, increasing the difficulty of accurately capturing it. Nevertheless, our framework learns graph distributions with a representation generator, which enables the modeling of complex underlying patterns. (3) Our framework GraphRCG is particularly competitive in the MMD score regarding degree distributions and the number of orbits. This observation demonstrates that GraphRCG could authentically capture the complex graph distribution of the training samples with the help of self-conditioned modeling. We include additional visualization results of graphs generated by our framework in Appendix D.
Table 2. Results of various methods on the QM9 Dataset.
Model Validity↑ NSPDK↓ FCD↓ Unique↑
EDP-GNN 47.52 0.005 2.68 99.25
SPECTRE 87.33 0.163 47.96 35.7
GDSS 95.72 0.003 2.9 98.46
DiGress 99.01 0.0005 0.36 96.66
GraphARM 90.25 0.002 1.22 95.62
GraphRCG 99.12 0.0008 0.28 98.39
Table 3. Results of various methods on the ZINC250k Dataset.
Model Validity↑ NSPDK↓ FCD↓ Unique↑
EDP-GNN 82.97 0.049 16.74 99.79
SPECTRE 90.20 0.109 18.44 67.05
GDSS 97.01 0.019 14.66 99.64
DiGress 91.02 0.082 23.06 81.23
GraphARM 88.23 0.055 16.26 99.46
GraphRCG 92.38 0.041 13.48 96.15
$\alpha=0$
(a) *
$\alpha=0.2$
(b) *
$\alpha=0.4$
(c) *
$\alpha=0.6$
(d) *
$\alpha=0.8$
(e) *
$\alpha=1.0$
(f) *
Figure 4. The generated graphs from the Planar dataset with different interpolation ratios between two representations.
4.3. Molecular Graph Generation
To evaluate our framework on molecular graph generation, we select two popular datasets: QM9 (Wu et al., 2018) and ZINC250k (Irwin et al., 2012). Particularly, QM9 consists of molecular graphs with a maximum of nine heavy atoms that could be treated with implicit or explicit hydrogens. We present the molecular graph generation results on QM9 in Table 2 and Table 3. Specifically, GraphRCG demonstrates competitive performance on QM9 across various metrics, compared to other state-of-the-art baselines. The best FCD values on QM9 and ZINC250k indicate that GraphRCG effectively captures the chemical property distributions in the dataset. Furthermore, the outstanding validity score on QM9 also signifies that our framework GraphRCG is capable of generating valid molecules that are more closely aligned with the training data distribution.
(a) SBM
(b) Planar
Figure 5. The t-SNE visualization of the Planar and SBM datasets, along with generated representations by our framework. Each point denotes the representation of a training graph sample or from our representation generator.
4.4. Representation Interpolation
Table 4. The ablation study results regarding self-conditioned modeling on dataset QM9 with explicit hydrogen.
Model Valid↑ Unique↑ Atom S.↑ Mol S.↑
Dataset 97.8 100 98.5 87.0
DiGress w/o A 92.3 97.9 97.3 66.8
DiGress w/ A 95.4 97.6 98.1 79.8
GraphRCG w/ A 92.9 95.4 93.1 76.6
GraphRCG w/ T 91.2 91.5 90.5 72.3
GraphRCG w/ T+M 94.6 97.0 90.2 74.9
GraphRCG w/ GMM 96.4 94.1 91.5 77.1
GraphRCG 96.9 98.1 97.2 81.9
4.5. Representation Quality
In this subsection, we provide visualization results to evaluate the quality of the generated representations. We present the t-SNE plot of representations from training samples and our representation generator on datasets SBM and Planar in Fig. 5. From the visualization results, we could observe that our representation generator could faithfully capture the graph distributions from training samples. In other words, the generated representations closely align with training graph distributions. More importantly, our framework can generate representations that slightly deviate from the training samples in specific directions. This indicates that our framework is capable of discovering novel graph distributions that are not present in training samples, while still ensuring the validity of these representations for further guidance during generations.
Table 5. Comparison of different denoising diffusion models for graph generation. $G(n,p)$ denotes the Erdős-Rényi graph model (Erdős et al., 1960), where $p$ is the probability of an edge existing between two nodes, and $n$ is the graph size.
Model Diffusion Type Convergence Conditional Generation Featured Graph Generation

### v2

@@ 3.2 Self-Conditioned Guidance (Training)
$(\widetilde{\mathbf{p}}^{\mathbf{X}},\widetilde{\mathbf{p}}^{\mathbf{E}})=g_{\theta}(\mathbf{X}_{t},\mathbf{E}_{t},\widetilde{\mathbf{h}}_{t},t),$ (7)
where $\widetilde{\mathbf{h}}_{t}=f_{\gamma}(\mathbf{h}_{t},t)$ is the denoised representation, generated from the representation generator at timestep $t$. Moreover, $\widetilde{\mathbf{p}}^{\mathbf{X}}\in\mathbb{R}^{a}$ and $\widetilde{\mathbf{p}}^{\mathbf{E}}\in\mathbb{R}^{b}$ are the predicted distributions for the node types and the edge types, respectively. The graph generator is based on the message-passing transformer architecture [44], as it effectively extracts the complex correlations between nodes and edges, while also being suitable for incorporating representations for conditioning [52]. Specifically, the layers incorporate the graph attention mechanism [51] into a Transformer framework [50], achieved by including normalization and feedforward layers. For each given (noisy) graph, the graph generator projects nodes and edges separately into low-dimensional representations and processes them through totally $L$ transformer layers, denoted as $(\mathbf{x}^{(l+1)}_{t},\mathbf{e}^{(l+1)}_{t})=M_{(l)}(\mathbf{x}^{(l+1)}_{t},\mathbf{e}^{(l+1)}_{t})$. Here $\mathbf{x}_{t}^{(l)}$ (or $\mathbf{e}_{t}^{(l)}$) is the representation for the nodes (or edges) in the $l$-th layer of the transformer, denoted as $M_{(l)}$.
Figure 3: The sampling process of our framework with self-conditioned guidance. Our step-wise incorporation strategy employs the denoised presentation at each timestep to guide the denoising of the noisy graph at the same time step, thereby progressively guiding each step closer to the learned distributions.
Cross-Attention. To effectively utilize the representations for guidance, we map them to the intermediate layers of the graph transformer via the cross-attention mechanism [50], achieved by
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}(QK^{T}/\sqrt{d})\cdot V.$ (8)
Specifically, the values of $Q$, $K$, and $V$ are computed as follows:
$Q=W_{Q}^{(l)}\cdot\mathbf{x}^{(l)},K=W_{K}^{(l)}\cdot\widetilde{\mathbf{h}}_{t},V=W_{V}^{(l+1)}\cdot\widetilde{\mathbf{h}}_{t},$ (9)
where $W_{Q}^{(l)}\in\mathbb{R}^{d\times d_{x}}$,$W_{K}^{(l)}\in\mathbb{R}^{d\times d_{h}}$, $W_{V}^{(l)}\in\mathbb{R}^{d\times d_{h}}$ are learneable projection matrices. $d_{x}$ and $d_{h}$ are the dimensions of $\mathbf{x}$ and $\mathbf{h}$, respectively. Notably, the above process is performed for edge representations $\mathbf{e}^{(l)}$ in the same way with different parameters.
@@
$\mathcal{L}_{GG}=\mathbb{E}_{G,t}\left[\sum\limits_{i=1}^{n}\text{CE}(\mathbf{X}_{i},\widetilde{\mathbf{p}}^{\mathbf{X}}_{i})+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\text{CE}(\mathbf{E}_{i,j},\widetilde{\mathbf{p}}^{\mathbf{E}}_{i,j})\right],$ (10)
@@
$\mathcal{L}_{AG}=\mathbb{E}_{G,t}\left[\left\|h_{\eta}(p^{\mathbf{X}},p^{\mathbf{E}})-f_{\gamma}\left(\mathbf{h}_{t},t\right)\right\|_{2}^{2}\right].$ (11)
@@
Note that $\mathcal{L}_{AG}$, together with $\mathcal{L}_{GG}$, will be only used for optimizing the graph generator, not involving the representation generator. In this way, we can ensure that the representation generator focuses on capturing the graph distribution. The detailed overall process of self-conditioned guidance is presented in Algorithm 2.
3.3 Self-Conditioned Guidance (Sampling)
$G_{t-1}\sim p_{\theta}(G_{t-1}|G_{t},\mathbf{h}_{t-1}),\ \ \text{where}\ \ \mathbf{h}_{t-1}=f_{\gamma}(\mathbf{h}_{t},t).$ (12)
In concrete, we keep track of the representation sampling process in the representation generator and utilize the representation in each step to guide the graph sampling in the same timestep $t$. In this manner, the generation process will absorb the graph distribution information learned by the representation generator, thereby improving generation performance.
4 Experiments
In our experiments, we evaluate GraphRCG across graph datasets covering realistic molecular and synthetic non-molecular datasets, in comparison to baselines, including auto-regressive models: GRAN [30] and GraphRNN [57], a GAN-based model: SPECTRE [33], and diffusion models: EDP-GNN [34], DiGress [52], GDSS [21], GraphARM [24], HiGen [22], and SparseDiff [36]. We provide implementation details and hyperparameter settings in Appendix D.2.
For generic graph generation, we evaluate the quality of the generated graphs with structure-based evaluation metrics. We follow previous work [57] and calculate the MMD (Maximum Mean Discrepancy) between the graphs in the test set and the generated graphs, regarding (1) degree distributions, (2) clustering coefficients distributions, (3) the number of orbits with four nodes, and (4) the spectra of the graphs obtained from the eigenvalues of the normalized graph Laplacian [57, 4]. For molecular graph generation, following previous works [21, 24], we evaluate the generated molecular graphs with several key metrics: (1) Frechet ChemNet Distance (FCD) [35], (2) Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD [5], (3) Validity, and (4) Uniqueness. We provide details of these metrics in Appendix D.3.
Table 1: Comparison of generation results on SBM, Planar, and Ego. The best results are shown in bold.
SBM Planar Ego
Training 0.0008 0.0332 0.0255 0.0063 0.0002 0.0310 0.0005 0.0052 0.0002 0.0100 0.0120 0.0014
SPECTRE 0.0015 0.0521 0.0412 0.0056 0.0005 0.0785 0.0012 0.0112 - - - -
HiGen 0.0019 0.0498 0.0352 0.0046 - - - - 0.0472 0.0031 0.0387 0.0062
4.1 Comparative Results
Generic Graph Generation. In this subsection, we further evaluate our framework on generic graph datasets with relatively larger sizes than molecular graphs. In particular, we consider two synthetic datasets: SBM, drawn from stochastic block models [33], with a maximum size of 200 nodes, and Planar, containing planar graphs with a fixed size of 64 [52]. In addition, we consider a realistic citation dataset Ego [42], originated from Citeseer [11]. Further details of these datasets are provided in Appendix D.1. From the results presented in Table 1, we could obtain the following observations: (1) GraphRCG outperforms other baselines on all three datasets across various metrics for graph generation, demonstrating the effectiveness of our framework in precisely capturing graph distributions and utilizing them for generation guidance. (2) The performance improvement over other methods is more substantial on the Planar dataset. As illustrated in the t-SNE plot in Fig. 7 (b), the distribution of the Planar dataset is more scattered, increasing the difficulty of accurately capturing it. Nevertheless, our framework learns graph distributions with a representation generator, which enables the modeling of complex underlying patterns. (3) GraphRCG is particularly competitive in the MMD score regarding degree distributions and the number of orbits. This observation demonstrates that GraphRCG could authentically capture the complex graph distribution of the training samples with the help of self-conditioned modeling. We include additional visualization results of graphs generated by our framework in Appendix F.
Table 2: Results of various methods on the QM9 and ZINC250k Datasets.
QM9 Dataset ZINC250k Dataset
Model Validity↑ NSPDK↓ FCD↓ Unique↑ Validity↑ NSPDK↓ FCD↓ Unique↑
EDP-GNN 47.52 0.005 2.68 99.25 82.97 0.049 16.74 99.79
SPECTRE 87.33 0.163 47.96 35.7 90.20 0.109 18.44 67.05
GDSS 95.72 0.003 2.9 98.46 97.01 0.019 14.66 99.64
DiGress 99.01 0.0005 0.36 96.66 91.02 0.082 23.06 81.23
GraphARM 90.25 0.002 1.22 95.62 88.23 0.055 16.26 99.46
GraphRCG 99.12 0.0008 0.28 98.39 92.38 0.041 13.48 96.15
Molecular Graph Generation. To evaluate our framework on molecular graph generation, we select two popular datasets: QM9 [54] and ZINC250k [19], with details and provided in Appendix D.1. We present the molecular graph generation results on QM9 in Table 2. Specifically, GraphRCG demonstrates competitive performance on QM9 across various metrics, compared to other state-of-the-art baselines. The best FCD values on QM9 and ZINC250k indicate that GraphRCG effectively captures the chemical property distributions in the dataset. Furthermore, the outstanding validity score on QM9 also signifies that our framework GraphRCG is capable of generating valid molecules that are more closely aligned with the training data distribution.
$\alpha=0.2$
(a) *
$\alpha=0.4$
(b) *
$\alpha=0.6$
(c) *
$\alpha=0.8$
(d) *
Figure 4: The generated graphs from Planar with different interpolation ratios between two representations.
4.2 Representation Interpolation
4.3 Ablation Study
Effect of Self-Conditioned Modeling. Previous works [52] have also explored various strategies to capture and model distributions for the guidance of graph generation. For example, DiGress leverages manually designed structural features (e.g., the number of cycles) for graph generation. However, such features contain less information and require domain knowledge. In this subsection, we compare several variants of our framework with these structural features to evaluate the efficacy of self-conditioned modeling in encapsulating the graph distribution. We consider using the following features (representations) to replace our self-conditioned modeling module: (1) structural features proposed in DiGress, (2) representations of true training samples, which limit the uniqueness of generation, (3) a mixture of true representations, and (4) distributions learned by a Gaussian Mixture Model (GMM). As the structural features involve molecular information, we conduct experiments on the QM9 dataset with explicit hydrogens, which is a more complex setting with larger graphs.
Table 3: The ablation study results regarding self-conditioned modeling on dataset QM9 with explicit hydrogen.
Model Valid↑ Unique↑ Atom S.↑ Mol S.↑
Dataset 97.8 100 98.5 87.0
DiGress w/o A 92.3 97.9 97.3 66.8
DiGress w/ A 95.4 97.6 98.1 79.8
GraphRCG w/ A 92.9 95.4 93.1 76.6
GraphRCG w/ T 91.2 91.5 90.5 72.3
GraphRCG w/ T+M 94.6 97.0 90.2 74.9
GraphRCG w/ GMM 96.4 94.1 91.5 77.1
GraphRCG 96.9 98.1 97.2 81.9
Table 4: The ablation study results regarding self-conditioned guidance on the dataset Ego.
Dataset Ego
GraphRCG-Gradient 0.0134 0.0955 0.0464 0.0151
GraphRCG w/o $\mathcal{L}_{AG}$ 0.0088 0.0647 0.0252 0.0092
GraphRCG-fixed 0.0053 0.0653 0.0310 0.0125
GraphRCG 0.0015 0.0448 0.0183 0.0042
Effect of Self-Conditioned Guidance. In this subsection, we investigate the effect of self-conditioned guidance in our framework. We first replace the entire guidance strategy with a gradient-based method, which leverages computed gradients as guidance. Therefore, the graph generator does not involve any representation during training. For the second variant, we remove the alignment loss $\mathcal{L}_{AG}$ described in Eq. (11). Without this loss, the graph generator is not well-aligned with the representations produced by the representation generator, thereby affecting the guidance performance. For the third variant, we directly use the fixed representation, i.e., the clean representation, to guide generation. In this case, the step-wise guidance strategy is removed, resulting in the infeasibility of progressive guidance. We provide further details regarding these variants in Appendix D.5. From the results presented in Table 4, we could first observe that our framework outperforms all other variants in most evaluation metrics, demonstrating the effectiveness of our self-conditioned guidance strategy. Moreover, GraphRCG-Gradient and GraphRCG w/o $\mathcal{L}_{AG}$ exhibit significantly poorer performance, as evidenced by higher scores across all metrics. This deterioration, particularly in Clus. and Orbit values, underscores the difficulty in preserving graph distributions without representation guidance or alignment loss. GraphRCG-fixed shows improved performance relative to the other variants, demonstrating the benefits of representation guidance even with a fixed one. Nevertheless, the results also indicate that the step-wise self-conditioned guidance is more beneficial for progressively guiding the generation process.
5 Conclusion
In this work, we investigate the importance of capturing and utilizing distributions for graph generation. We propose a novel self-conditioned generation, that encompasses self-conditioned modeling and self-conditioned guidance. Instead of directly learning from graph distributions, we encode all graphs into representations to capture graph distributions, which could preserve richer information. Our self-conditioned guidance module further guides the generation process in each timestep with representations with different degrees of noise. We conduct extensive experiments to evaluate our framework, and the results demonstrate the efficacy of our framework in graph generation.
Appendix A Noise Model
$\mathbf{Q}_{t}^{\mathbf{X}}=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{a}(\mathbf{p}^{\mathbf{X}})^{\top},\ \ \text{and}\ \ \mathbf{Q}_{t}^{\mathbf{E}}$ $=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{b}(\mathbf{p}^{\mathbf{E}})^{\top}.$ (13)
In practice, the noise process is not cumulative. By defining $\overline{\alpha}^{t}=\prod_{i=1}^{t}\alpha^{i}$ and $\overline{\beta}^{t}=1-\overline{\alpha}^{t}$, we could achieve
$\prod_{i=1}^{t}\mathbf{Q}_{i}^{\mathbf{X}}=\overline{\alpha}^{t}\mathbf{I}+\overline{\beta}^{t}\mathbf{1}_{a}(\mathbf{p}^{\mathbf{X}})^{\top},\ \ \text{and}\ \ \prod_{i=1}^{t}\mathbf{Q}_{i}^{\mathbf{E}}=\overline{\alpha}^{t}\mathbf{I}+\overline{\beta}^{t}\mathbf{1}_{a}(\mathbf{p}^{\mathbf{E}})^{\top}.$ (14)
Following [52], we set $\overline{\alpha}^{t}=\cos(0.5\pi(t/T+s)/(1+s))^{2}$ with $s$ being a small value. In this manner, when computing the noisy graphs $G_{t}=(\mathbf{X}_{t},\mathbf{E}_{t})$, we do not need to recursively multiply $Q_{t}^{\mathbf{X}}$. Instead, we utilize the value of $\overline{\alpha}^{t}$ to directly obtain $G_{t}$.
Appendix B Self-conditioned Modeling
Algorithm 1 The training process of self-conditioned modeling.
0:  Optimized graph encoder and representation generator.
1:  for $i=1,2,\dotsc,T_{tr}$ do
2:     Sample $G=(\mathbf{X},\mathbf{E})$ from $\mathcal{G}$;
4:     $\mathbf{h}_{0}\leftarrow h_{\eta}(\mathbf{X},\mathbf{E})$;
6:     $\widetilde{\mathbf{h}}_{t}=f_{\gamma}\left(\mathbf{h}_{t},t\right)$;
9:     $\overline{\mathbf{h}}_{t}=h_{\eta}(\mathbf{X}_{t},\mathbf{E}_{t})$;
10:     Optimize $h_{\eta}$ with $\mathcal{L}_{AR}$ in Eq. (3);
11:  end for
Appendix C Self-conditioned Guidance
@@
C.1 Comparison with Other Methods
In this subsection, we discuss and compare our framework GraphRCG with other diffusion-based generative models for graph data. We provide the comparisons regarding technical details in Table 5. Typically, existing diffusing methods for graph generation rely on either continuous or discrete diffusion. As a classific example, GDSS [21] adopts Gaussian transition kernels to perform continuous diffusion with a score-matching strategy. DiGress, on the other hand, performs discrete diffusion while considering categorical features of nodes and edges. Nevertheless, although continuous representations could capture complex structural patterns in graph distributions, they are less effective in generating discrete graph data. In contrast, our framework is capable of leveraging continuous diffusion to guide discrete diffusion, thereby combining the strengths of both and allowing for the generation of a wider range of graph structures. Regarding the convergence type, which refers to the pure noise state, we also combine both continuous and discrete noise to facilitate the optimization of both our representation generator and graph generator. Comparing the conditional generation type, our framework is conditioned on representations, which preserve richer information of complex distributions of each dataset. As a result, the generation process could significantly benefit from the guidance of representations.
Table 5: Comparison of different denoising diffusion models for graph generation. $G(n,p)$ denotes the Erdős-Rényi graph model [9], where $p$ is the probability of an edge existing between two nodes, and $n$ is the graph size.
Model Diffusion Type Convergence Conditional Generation Featured Generation